UT 20161019 LUCI1 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt
Observer AO: DMiller
Telescope Operator: GBechetti

Plan for the night

- If seeing is good: - check LUCI focus & NCPA and than do LUCI1 AO-com

- if not, ARGOS-team starts and we may interrupt later tonight

Summary

- mostly bugfixing and/or finding them

- 3 Bayside stage errors

- within a script big offset off the patrol fieldis possible but does'nt go back, this bug must be removed

- for very bright AOref stars a ND-filter must be put in FLAO, implemented (and hopefully working reliably)

Details

All times are UT
Files are luci1.20161019.NNNN.fits

00:50 Sunset, Weather: photometric, T=9deg, Hum=30%, Wind 5-10 m/s

The ARGOS-team is starting with twilight flats

01:20 ARGOS is done with flats, we take over, need to wait a bit as Geno's having
problems with the DX-ADSec

01:35 Geno checks pointing and collimates

01:43 Seeing looks good, guider reports about 0.7"

LUCI1-AO robustness & Stability

01:45 Preset to P Cygni, doesn't work, preset failed, get stuck
LUCI apparently doesn't talk to IIFDoug is doing a preset via IDL, pointing off,
check & correct pointing, redo preset to PCyg, OK, RunAO failed, Bayside stage error
resent preset, RunAO, failed again, home problem, resent preset once more, failed again, hotspot saturated, star is too bright!!!

02:25 Doug changes config-file to dimm the star, now 90% of light goes to acq-camera and
only 10% to the Pyramid, UAO thinks now its a 9th mag star

02:27 4th attempt to grab PCyg, Bayside stage error, resent preset,
RunAO worked, 1st image has 5pix FWHM, 1st big 600" offset OK, 2nd big offset failed,
more than one big offset outside the patrol field is apparently not in the logic,
bug, large offsets were possible in March/April when we observed M5 and must be possible ,
and resent preset again, RunAO, failed (ich glaubs nicht!!!), hotspot still saturated,
Doug put in an even stronger filter, resent preset, timeout filter wheel 1 (at FLAO),
resent preset, RunAO Ok, loop opened during 1st ima and closed back safe skip

Stop working on that, too many bugs to be removed

images #01xx-0119

03:21 Preset to 3C 452, has offsets which are within the patrol field, small ones and big ones
RunAO failed, ND-filter was not put out automatically (although it should), resent preset, RunAO,
bin=2 (AOref has 10.8 in nomad and 12.8 in FLAO), loop opened after 23 images,
human interaction communication error when changing gain, loop manually, closed,
script resettet and skiped to the relevant item and cont'd, OK, than skip frame,
shell RIPed, #0143 is crap, need to resent presest again, skip frame once more,
resent preset, failed

images #0119-0153

04:57 End of tests

dimm.20161018jhill.jpg

-- %USERSIG{JochenHeidt - 2016-10-18}%
Topic revision: r10 - 07 Nov 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback